OpenStudy (anonymous):

Hello all, I'm having trouble with a few problems on my pretest for Calculus II, so here goes the first: Evaluate the integral using the indicated trigonometric substitution Integrate[1/(x^2sqrt(x^2-16)) , {x}]; x=4secant theta

OpenStudy (2bornot2b):

I think you would have got better answers or rather you will get more users to answer your question, if you would have posted the question alone, I mean try excluding the part "Hello all...........here goes the first"

OpenStudy (anonymous):

Oh, mkay, I'll try that. Thanks!

OpenStudy (2bornot2b):

And you can also try posting the question with the equation in LaTeX format, in that way you can increase the probability too.

OpenStudy (anonymous):

Latex?

OpenStudy (2bornot2b):

Do you want me to help you out, by posting it on behalf of you?

OpenStudy (anonymous):

Yeah, that'd be awesome!

OpenStudy (anonymous):

Yeah, with this before it: Evaluate the integral using the indicated trigonometric substitution

OpenStudy (anonymous):

and no ,x

OpenStudy (2bornot2b):

And I think it must be dx instead of that x right?

OpenStudy (anonymous):

its dx*

OpenStudy (anonymous):

yeah lol

OpenStudy (2bornot2b):

Ok now see this one "Evaluate the integral \[\int\frac{1}{(x^2\sqrt{x^2-16)}}dx\] using the substitution \[x=4\sec\theta\]" Is it OK?

OpenStudy (anonymous):

yes

OpenStudy (2bornot2b):

Ok so I am posting it, ...

OpenStudy (anonymous):

Mkay, thanks

OpenStudy (2bornot2b):

OK, so its posted at http://openstudy.com/study#/updates/4f2c150ee4b0e1bedb5d1f0a

OpenStudy (mr.math):

Good tips. Questions written with Latex always attract me :-D

OpenStudy (2bornot2b):

And I also used the chat rooms to invite people in here. And I messaged myin using fan message

OpenStudy (2bornot2b):

So I think I can summarize them as follows 1. Make it concise 2. Use equation editor 3. Use the chatrooms to attract users 4. Private message the people you think can solve the problem

OpenStudy (anonymous):

I didn't know this place had pm's 0-0